OpenStudy (anonymous):

I NEED VECTOR HELP!!! CAN ANYONE HERE DO CAL 2 CALCULATIONS???

OpenStudy (anonymous):

Find two unit vectors that make an angle of 60° with v = ‹6, 8›. Give your answers correct to three decimal places.

OpenStudy (anonymous):

ANYONE????

OpenStudy (anonymous):

I tried, sitting this one out. Will be watching with great interest.

OpenStudy (anonymous):

ROACHIE sorry :(

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

ill give you a medal if someone figures this out...lol

OpenStudy (anonymous):

ah!

OpenStudy (anonymous):

what is the magnitude of u?

OpenStudy (anonymous):

|v|=10

OpenStudy (anonymous):

6a + 8b = \[\left| u \right|\] (10)(1/2)

OpenStudy (anonymous):

... u?

OpenStudy (anonymous):

by theorem 3 we need u * v = \[\left| v \right| \left| u \right|\] cos 60

OpenStudy (anonymous):

would u be 1?

OpenStudy (anonymous):

\[\sqrt{a ^{2}+b^{2}}\]

OpenStudy (anonymous):

if i knew what the magnitude of u was i could figure this problem out.

OpenStudy (anonymous):

oh, u=<6,8>??

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

u = <a,b>

OpenStudy (anonymous):

v = <6,8>

OpenStudy (dumbcow):

the magnitude of u (unit vector) is 1 i dont remember the formula for that theorem you were referring

OpenStudy (anonymous):

yea, I dont know where u came from but yes, u being a unit vector has magnitude 1

OpenStudy (anonymous):

6a + 8b = 5 therefore b=5/8 - 3/4a

OpenStudy (anonymous):

a^2 + (5/8 -3/4a)^2 = 1

OpenStudy (anonymous):

?? any thoughts on that?

OpenStudy (dumbcow):

where does the 5 come from?

OpenStudy (anonymous):

1/2 (10)

OpenStudy (anonymous):

u * v = |v||u| cos 60 now I see u, duh

OpenStudy (dumbcow):

why are we dividing 10 (|v|) by 2

OpenStudy (anonymous):

cos of 60 is 1/2

OpenStudy (anonymous):

someone want to factor that out for me...

OpenStudy (anonymous):

because im sure to f THAT up.

OpenStudy (anonymous):

if we know that u is a unit vector then v * u = |v|cos 60 if you unit vector v then v*u = cos 60 that how I think of it...

OpenStudy (anonymous):

im not following r . . . . .

OpenStudy (dumbcow):

haha...25/64 -30/32a +9/16a^2

OpenStudy (anonymous):

v-hat = v/|v|

OpenStudy (anonymous):

then quadratic... fun fun

OpenStudy (anonymous):

|v-hat|=1

OpenStudy (anonymous):

30/32 +- \[\sqrt{30/32^{2}-4(25/64)(9/16)}/2(25/64)\]

OpenStudy (anonymous):

fml

OpenStudy (anonymous):

did you add the original a^2 to the factor? dumbcow?

OpenStudy (dumbcow):

no i didn;t multiply everything by 64 100a^2 -60a -39 = 0 a = -.3928, .9928

OpenStudy (anonymous):

then i need to plug those back into 5/8 -3/4a to find b right?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

<-.3928,.9196> and <.9928, -.1196>

OpenStudy (dumbcow):

looks good

OpenStudy (anonymous):

YES!!!!!

OpenStudy (anonymous):

thank god!